# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ IV.4.2 Convergence of conformal structure
Hence the condition (IV.62b) holds for $j=j_{0}+1$. Since $\xi_{j_{0}}$ vanishes on a neighborhood of $\Sigma\setminus\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}$, and $\omega_{k,j_{0}}\circ\varphi_{\infty,j_{0}+1}\in C^{\infty}\big(\big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)\cap U_{j_{0}+1}\big)$ we have $(\xi_{j_{0}}\circ\varphi^{-1}_{\infty,j_{0}+1})\,\omega_{k,j_{0}}\in C^{\infty}(\varphi_{\infty,j_{0}+1}(U_{j_{0}+1}))$. Then we define
@@
Since we have $\overline{V_{j_{0}+1,j_{0}+1}}=\overline{U^{\prime}_{j_{0}+1}}\subset U_{j_{0}+1}$, the condition (IV.62d) is thus satisfied for $\beta=j=j_{0}+1$, $\alpha\leq j_{0}$. By using (IV.62a) and the identity $\varphi_{k,\alpha,\beta}=\varphi_{k,\beta,\alpha}^{-1}$, we then obtain (IV.62d) for $j=j_{0}+1$ and any $\alpha,\beta\leq j_{0}+1$.
Next we prove $\rho_{k,j_{0}+1}\rightarrow id$ in $C^{\infty}(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}}),\mathbb{R}^{2})$ as $k\rightarrow\infty$. Fixing $\alpha\leq j_{0}$, let $W$ be an open set such that $\varphi_{\infty,\alpha}(\overline{V_{\alpha,j_{0}}}\cap\overline{U_{j_{0}+1}^{\prime}})\subset W\Subset\varphi_{\infty,\alpha}(U_{\alpha}\cap U_{j_{0}+1})$. The convergence (IV.60) implies that
@@
By our induction hypothesis (IV.62c) for $\alpha\leq j_{0}$, we also have
@@
$\omega_{k,j_{0}}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}(\varphi_{\infty,j_{0}+1}(\overline{V_{\alpha,j_{0}}}\cap\overline{U^{\prime}_{j_{0}+1}})$.}$
@@
$\rho_{k,j_{0}+1}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}\Big(\varphi_{\infty,j_{0}+1}\Big(\Big(\bigcup_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\Big)\cap\overline{U^{\prime}_{j_{0}+1}}\Big)\Big)$.}$
@@
Hence the condition (IV.62c) holds for $j=j_{0}+1$. In particular, $\rho_{k,j_{0}+1}$ is an immersion on $\varphi_{\infty,j_{0}+1}(\overline{V_{j_{0}+1,j_{0}+1}})=\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})$ for $k$ large enough. Also, since $\partial U^{\prime}_{j_{0}+1}$ is smooth, we have the following embedding (see [58, Theorem 5.8.4]):
$W^{1,\infty}(\varphi_{\infty,j_{0}+1}(U^{\prime}_{j_{0}+1}))\subset C^{0,1}(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})).$
@@
Now since $\Sigma=\bigcup\limits_{1\leq\alpha\leq l}V_{\alpha}$, we can choose smooth non-negative cut-off functions $\phi_{\alpha}\in C^{\infty}_{c}(\varphi_{\infty,\alpha}(V_{\alpha}))$ such that $\sum\limits_{\alpha=1}^{l}\phi_{\alpha}\circ\varphi_{\infty,\alpha}>0$ on $\Sigma$. Since $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$ in $C^{0}(\overline{V_{\alpha}})$ as $k\rightarrow\infty$, we have $\phi_{\alpha}\in C_{c}^{\infty}(\varphi_{k,\alpha}(V_{\alpha}))$ for $k\in\mathbb{N}\cup\{\infty\}$ large enough. For such $k$ we define
@@
Then $g_{k}^{\prime}$ is a smooth metric under the complex structure induced by $\vec{\Phi}_{k}$ as defined in Corollary IV.5. Since $\varphi^{*}_{k,\alpha}\,g_{\text{std}}$ is weakly conformal to $g_{k}$ on $V_{\alpha}$, we have $g_{k}^{\prime}$ is weakly conformal to $g_{k}$. Since we assumed $\tilde{g}$ and $g_{\infty}$ are weakly conformal, we obtain that $\varphi_{\infty,\alpha}$ is weakly conformal from $(V_{\alpha},g)$ to $(\varphi_{\infty,\alpha}(V_{\alpha}),g_{\text{std}})$. Hence $\varphi_{\infty,\alpha}$ is smooth by Corollary IV.5 (i). Let $1\leq\alpha\leq l$, $x\in V_{\alpha}$, then by (IV.60) and (IV.61) we have
@@
Assume $\tilde{g_{k}}\rightarrow\tilde{g}_{\infty}$ a.e. Then we can construct the conformal charts $\{\varphi_{k,\alpha}\}$ as in Remark IV.6, and by using the same argument as in Theorem IV.19, we find $\{h_{k}\}$, $\{\Psi_{k}\}$ satisfying Theorem IV.19 (i), (ii), (iv), but $\Psi_{k}$ may not be Lipschitz. Indeed, by [90, Proposition 4.36] and [91], we have $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$, $\varphi^{-1}_{k,\alpha}\rightarrow\varphi^{-1}_{\infty,\alpha}$ in $W^{1,p}_{\textup{loc}}$ for some $p>2$, hence $\Psi_{k}\rightarrow id$, $\Psi^{-1}_{k}\rightarrow id$ in $W^{1,p}(\Sigma,\Sigma)$, in particular, in $C^{0}(\Sigma,\Sigma)$.
@@ IV.4.3 Approximation results
The first goal of this section is to prove Theorem IV.23, namely that a $W^{1,\infty}$-immersion can be approximated by smooth immersions as soon as its first derivative belongs to $VMO$. Then we prove Theorem I.3 at the end. We define a weak $W^{1,\infty}$-immersion with value into manifolds.
@@
We estimate the mean oscillation of a composition and the distance with a smooth approximation.
@@
$\beta_{r}(\nabla(f\circ\varphi^{-1}))\leq C\left(\beta_{r}(\nabla f)+\beta_{r}(\nabla\varphi)\|\nabla f\|_{L^{\infty}(U)}\right).$ (IV.63)
@@
$\begin{aligned} &\left|\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))-\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})\right|\\[2.84526pt] &\leq C\Big(\beta_{\varepsilon}(\nabla f)+\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\|\nabla f\|_{L^{\infty}(U)}\Big),\end{aligned}$ (IV.64)
@@
The estimate (IV.63) is thus proved.
@@
Hence it holds that $|x_{0}-\varphi^{-1}(y)|\leq\varepsilon\,(K+\|\nabla\varphi^{-1}\|_{L^{\infty}(W)})$. Note that $(\nabla\varphi)^{-1}\in C^{0}(\overline{U})$ and for all $r>0$, we have
@@
$\begin{aligned} &\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})-\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}\right|\\[5.69054pt] &\leq C\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\,\|\nabla f\|_{L^{\infty}(U)}.\end{aligned}$ (IV.65)
Since $\operatorname{dist}(\varphi(x_{1}),\partial W)>\varepsilon$, we have $\operatorname{dist}(x_{1},\partial U)>\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}(U)}$. Now we can apply the proof of Riemann’s Theorem [162, Theorem 2] and Lemma III.13 to get
$\bigg|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\nabla f).$ (IV.66)
@@
$\begin{aligned} \bigg|\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\tilde{f})\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.67)
@@
$\begin{aligned} \left|\nabla f_{\varepsilon}(x_{0})-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|&=\left|\int_{B_{\varepsilon}(x_{0})}\eta_{\varepsilon}(x_{0}-x)\,\left(\nabla f(x)-\fint_{B_{\varepsilon}(x_{0})}\nabla f\,dx\right)\right|\\[2.84526pt] &\leq C\,\fint_{B_{\varepsilon}(x_{0})}\left|\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|\\[2.84526pt] &\leq C\,\beta_{\varepsilon}(\nabla f).\end{aligned}$ (IV.68)
Combining (IV.66), (IV.67), (IV.68) we obtain by the triangle inequality
$\begin{aligned} &\left|\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))\right|\\[2.84526pt] &\leq\|\nabla\varphi^{-1}\|_{L^{\infty}}\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla f_{\varepsilon}(x_{0})\right|\\[2.84526pt] &\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.69)
@@
We are now ready to establish an approximation result required for the proof of Theorem I.3. The underlying idea goes back to [178], see also [100, Remark 2.1]. For completeness, we present a detailed proof.
@@
$\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{k})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}},\qquad\text{for all }\,v\in TU\,\text{ and all }\,k\in\mathbb{N}.$ (IV.70)
@@
Since $\partial U$ is Lipschitz and $\in W^{1,\infty}(U), ff$ has a unique Lipschitz continuous representative defined on $\overline{U}$. In particular, $f(\overline{U})$ is compact. Since $M$ is smooth, there exists a bounded tubular neighborhood $W$ of $M$ with a retraction map $\pi\in C^{\infty}(W,M)$ such that $\pi(y)=y$ for any $y\in M$. Since $f(U)$ is compact, there exists a precompact neighborhood $W^{\prime}$ of $f(U)$ strictly contained in $W$. Since $\overline{U}$ is compact, there exist coordinate systems $\{(U_{j},\phi_{j})\}_{j=1}^{s}$ of $N$ such that $\overline{U}\subset\bigcup_{j=1}^{s}U_{j}$. Since $\partial U$ is Lipschitz, we can further assume that for each $U_{j}\nsubset U$, we have $\phi_{j}(U_{j})=B^{n}\subset\mathbb{R}^{n}$, $0\in\phi_{j}(\partial U)$ and, for some Lipschitz map $\gamma_{j}\colon\mathbb{R}^{n-1}\rightarrow\mathbb{R}$,
@@
$0<\varepsilon<\min_{1\leq j\leq s}\left[\frac{1}{K_{j}+1}\min\Big(\text{dist}\big(\phi_{j}(\operatorname{supp}(\xi_{j})),\phi_{j}(\partial V_{j})\big),\text{dist}\big(\phi_{j}(V_{j}),\phi_{j}(U_{j})^{c}\big)\Big)\right].$ (IV.71)
@@
By choice of $\varepsilon$ in (IV.71), it holds that
@@
$\forall v\in T\Sigma,\qquad\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{\varepsilon})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}}$. (IV.72)
@@
$\begin{aligned} &\left|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(f\circ\phi_{j}^{-1})\right|\\ &\leq\sum_{j^{\prime}=1}^{s}\left|\nabla(F_{j^{\prime},\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(\tilde{f}_{j^{\prime}}\circ\phi_{j}^{-1})\right|\\ &=\underset{\varepsilon\to 0}{o}(1).\end{aligned}$ (IV.73)
@@
$\fint_{B_{\varepsilon}(z^{\varepsilon})}\left|\nabla(f\circ\phi_{j}^{-1})-\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)\right|=\underset{\varepsilon\to 0}{o}(1).$ (IV.74)
@@
$\alpha\coloneqq\|\nabla\pi\|_{C^{0}(\overline{W^{\prime}})}+1+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}.$ (IV.76)
By (IV.74), there exists $\varepsilon_{0}>0$ such that for any $\varepsilon\in(0,\varepsilon_{0})$, we have on the one hand that $F_{\varepsilon}\in C^{\infty}(\overline{U},M)$ is well-defined with $F_{\varepsilon}(\overline{U})\subset W^{\prime}$. On the other hand for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, there holds $B_{\varepsilon}(z^{\varepsilon})\subset\phi_{j}(V_{j}\cap\overline{U})$, and there exists $z^{\prime}\in B_{\varepsilon}(z^{\varepsilon})$ satisfying $f\circ\phi_{j}^{-1}$ is differentiable at $z^{\prime}$ with the following inequalities:
@@
    $\big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\leq(2\alpha)^{-1}C_{j}.$ (IV.77)
@@
    $\begin{aligned} &\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z)\big|+\big|(f\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq(2\alpha)^{-1}C_{j}.\end{aligned}$ (IV.78)
@@
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)\big|\\[5.69054pt] &\geq\ \big|d(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|-\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|.\end{aligned}$ (IV.79)

### v3

@@ IV.4.2 Convergence of conformal structures
Hence the condition (IV.63b) holds for $j=j_{0}+1$. Since $\xi_{j_{0}}$ vanishes on a neighborhood of $\Sigma\setminus\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}$, and $\omega_{k,j_{0}}\circ\varphi_{\infty,j_{0}+1}\in C^{\infty}\big(\big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)\cap U_{j_{0}+1}\big)$ we have $(\xi_{j_{0}}\circ\varphi^{-1}_{\infty,j_{0}+1})\,\omega_{k,j_{0}}\in C^{\infty}(\varphi_{\infty,j_{0}+1}(U_{j_{0}+1}))$. Then we define
@@
Since we have $\overline{V_{j_{0}+1,j_{0}+1}}=\overline{U^{\prime}_{j_{0}+1}}\subset U_{j_{0}+1}$, the condition (IV.63d) is thus satisfied for $\beta=j=j_{0}+1$, $\alpha\leq j_{0}$. By using (IV.63a) and the identity $\varphi_{k,\alpha,\beta}=\varphi_{k,\beta,\alpha}^{-1}$, we then obtain (IV.63d) for $j=j_{0}+1$ and any $\alpha,\beta\leq j_{0}+1$.
Next we prove $\rho_{k,j_{0}+1}\rightarrow id$ in $C^{\infty}(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}}),\mathbb{R}^{2})$ as $k\rightarrow\infty$. Fixing $\alpha\leq j_{0}$, let $W$ be an open set such that $\varphi_{\infty,\alpha}(\overline{V_{\alpha,j_{0}}}\cap\overline{U_{j_{0}+1}^{\prime}})\subset W\Subset\varphi_{\infty,\alpha}(U_{\alpha}\cap U_{j_{0}+1})$. The convergence (IV.61) implies that
@@
By our induction hypothesis (IV.63c) for $\alpha\leq j_{0}$, we also have
@@
$\omega_{k,j_{0}}\xrightarrow[k\rightarrow\infty]{}id\qquad\text{in $C^{\infty}\left(\varphi_{\infty,j_{0}+1}(\overline{V_{\alpha,j_{0}}}\cap\overline{U^{\prime}_{j_{0}+1}})\right)$.}$
@@
$\rho_{k,j_{0}+1}\xrightarrow[k\rightarrow\infty]{}id\qquad\text{in $C^{\infty}\Big(\varphi_{\infty,j_{0}+1}\Big(\Big(\bigcup_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\Big)\cap\overline{U^{\prime}_{j_{0}+1}}\Big)\Big)$.}$
@@
Hence the condition (IV.63c) holds for $j=j_{0}+1$. In particular, $\rho_{k,j_{0}+1}$ is an immersion on $\varphi_{\infty,j_{0}+1}(\overline{V_{j_{0}+1,j_{0}+1}})=\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})$ for $k$ large enough. Also, since $\partial U^{\prime}_{j_{0}+1}$ is smooth, we have the following embedding (see [58, Theorem 5.8.4]):
$W^{1,\infty}\left(\varphi_{\infty,j_{0}+1}(U^{\prime}_{j_{0}+1})\right)\subset C^{0,1}\left(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})\right).$
@@
Since $\Sigma=\bigcup\limits_{1\leq\alpha\leq l}V_{\alpha}$, we can choose smooth non-negative cut-off functions $\phi_{\alpha}\in C^{\infty}_{c}(\varphi_{\infty,\alpha}(V_{\alpha}))$ such that $\sum\limits_{\alpha=1}^{l}\phi_{\alpha}\circ\varphi_{\infty,\alpha}>0$ on $\Sigma$. Since $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$ in $C^{0}(\overline{V_{\alpha}})$ as $k\rightarrow\infty$, we have $\phi_{\alpha}\in C_{c}^{\infty}(\varphi_{k,\alpha}(V_{\alpha}))$ for $k\in\mathbb{N}\cup\{\infty\}$ large enough. For such $k$ we define
@@
Then $g_{k}^{\prime}$ is a smooth metric under the complex structure induced by $\vec{\Phi}_{k}$ as defined in Corollary IV.5. Since $\varphi^{*}_{k,\alpha}\,g_{\text{std}}$ is weakly conformal to $g_{k}$ on $V_{\alpha}$, we have $g_{k}^{\prime}$ is weakly conformal to $g_{k}$. Since we assumed $\tilde{g}$ and $g_{\infty}$ are weakly conformal, we obtain that $\varphi_{\infty,\alpha}$ is weakly conformal from $(V_{\alpha},g)$ to $(\varphi_{\infty,\alpha}(V_{\alpha}),g_{\text{std}})$. Hence $\varphi_{\infty,\alpha}$ is smooth by Corollary IV.5 (i). Let $1\leq\alpha\leq l$, $x\in V_{\alpha}$, then by (IV.61) and (IV.62) we have
@@
Assume $\tilde{g_{k}}\rightarrow\tilde{g}_{\infty}$ a.e. Then we can construct the conformal charts $\{\varphi_{k,\alpha}\}$ as in Remark IV.6, and by using the same argument as in Theorem IV.19, we find $\{h_{k}\}$, $\{\Psi_{k}\}$ satisfying Theorem IV.19 (i), (ii), (iv), but $\Psi_{k}$ may not be Lipschitz. Indeed, by [91, Proposition 4.36] and [92], we have $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$, $\varphi^{-1}_{k,\alpha}\rightarrow\varphi^{-1}_{\infty,\alpha}$ in $W^{1,p}_{\textup{loc}}$ for some $p>2$, hence $\Psi_{k}\rightarrow id$, $\Psi^{-1}_{k}\rightarrow id$ in $W^{1,p}(\Sigma,\Sigma)$, in particular, in $C^{0}(\Sigma,\Sigma)$.
@@ IV.4.3 Approximation results
The first goal of this section is to prove Theorem IV.23, namely that a $W^{1,\infty}$-immersion can be approximated by smooth immersions as soon as its first derivative belongs to $VMO$. Then we prove Theorem I.3 at the end. We define a weak $W^{1,\infty}$-immersion taking values in manifolds.
@@
We estimate the mean oscillation of a composition and the distance to a smooth approximation.
@@
$\beta_{r}(\nabla(f\circ\varphi^{-1}))\leq C\left(\beta_{r}(\nabla f)+\beta_{r}(\nabla\varphi)\|\nabla f\|_{L^{\infty}(U)}\right).$ (IV.64)
@@
$\begin{aligned} &\left|\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))-\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})\right|\\[5.69054pt] &\leq C\Big(\beta_{\varepsilon}(\nabla f)+\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\|\nabla f\|_{L^{\infty}(U)}\Big),\end{aligned}$ (IV.65)
@@
The estimate (IV.64) is thus proved.
@@
Hence it holds that $|x_{0}-\varphi^{-1}(y)|\leq\varepsilon\,(K+\|\nabla\varphi^{-1}\|_{L^{\infty}(W)})$. Moreover it holds $(\nabla\varphi)^{-1}\in C^{0}(\overline{U})$ and for all $r>0$, we have
@@
$\begin{aligned} &\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})-\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}\right|\\[5.69054pt] &\leq C\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\,\|\nabla f\|_{L^{\infty}(U)}.\end{aligned}$ (IV.66)
Since $\operatorname{dist}(\varphi(x_{1}),\partial W)>\varepsilon$, we have $\operatorname{dist}(x_{1},\partial U)>\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}(U)}$. Now we can apply the proof of Riemann’s Theorem [165, Theorem 2] and Lemma III.13 to get
$\bigg|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\nabla f).$ (IV.67)
@@
$\begin{aligned} \bigg|\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\tilde{f})\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.68)
@@
$\begin{aligned} \left|\nabla f_{\varepsilon}(x_{0})-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|&=\left|\int_{B_{\varepsilon}(x_{0})}\eta_{\varepsilon}(x_{0}-x)\,\left(\nabla f(x)-\fint_{B_{\varepsilon}(x_{0})}\nabla f\,dx\right)\right|\\[2.84526pt] &\leq C\,\fint_{B_{\varepsilon}(x_{0})}\left|\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|\\[2.84526pt] &\leq C\,\beta_{\varepsilon}(\nabla f).\end{aligned}$ (IV.69)
Combining (IV.67), (IV.68), (IV.69) we obtain by the triangle inequality
$\begin{aligned} &\left|\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))\right|\\[2.84526pt] &\leq\|\nabla\varphi^{-1}\|_{L^{\infty}}\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla f_{\varepsilon}(x_{0})\right|\\[2.84526pt] &\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.70)
@@
We are now ready to establish an approximation result required for the proof of Theorem I.3. The underlying idea goes back to [181], see also [103, Remark 2.1] for the $W^{2,2}$ case in dimension $2$. For completeness, we present a detailed proof.
@@
$\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{k})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}},\qquad\text{for all }\,v\in TU\,\text{ and all }\,k\in\mathbb{N}.$ (IV.71)
@@
Since $\partial U$ is Lipschitz and $\in W^{1,\infty}(U), ff$ has a unique Lipschitz continuous representative defined on $\overline{U}$. In particular, $f(\overline{U})$ is compact. Since $M$ is smooth, there exists a bounded tubular neighborhood $W$ of $M$ with a retraction map $\pi\in C^{\infty}(W,M)$ such that $\pi(y)=y$ for any $y\in M$. Since $f(U)$ is compact, there exists a precompact neighborhood $W^{\prime}$ of $f(U)$ strictly contained in $W$. Since $\overline{U}$ is compact, there exist finitely many coordinate charts $\{(U_{j},\phi_{j})\}_{j=1}^{s}$ of $N$ such that $\overline{U}\subset\bigcup_{j=1}^{s}U_{j}$. Since $\partial U$ is Lipschitz, we can further assume that for each $U_{j}\nsubset U$, we have $\phi_{j}(U_{j})=B^{n}\subset\mathbb{R}^{n}$, $0\in\phi_{j}(\partial U)$ and, for some Lipschitz map $\gamma_{j}\colon\mathbb{R}^{n-1}\rightarrow\mathbb{R}$,
@@
$0<\varepsilon<\min_{1\leq j\leq s}\left[\frac{1}{K_{j}+1}\min\Big(\text{dist}\big(\phi_{j}(\operatorname{supp}(\xi_{j})),\phi_{j}(\partial V_{j})\big),\text{dist}\big(\phi_{j}(V_{j}),\phi_{j}(U_{j})^{c}\big)\Big)\right].$ (IV.72)
@@
By choice of $\varepsilon$ in (IV.72), it holds that
@@
$\forall v\in T\Sigma,\qquad\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{\varepsilon})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}}$. (IV.73)
@@
$\begin{aligned} &\left|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(f\circ\phi_{j}^{-1})\right|\\ &\leq\sum_{j^{\prime}=1}^{s}\left|\nabla(F_{j^{\prime},\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(\tilde{f}_{j^{\prime}}\circ\phi_{j}^{-1})\right|\\ &=\underset{\varepsilon\to 0}{o}(1).\end{aligned}$ (IV.74)
@@
$\fint_{B_{\varepsilon}(z^{\varepsilon})}\left|\nabla(f\circ\phi_{j}^{-1})-\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)\right|=\underset{\varepsilon\to 0}{o}(1).$ (IV.75)
@@
$\alpha\coloneqq\|\nabla\pi\|_{C^{0}(\overline{W^{\prime}})}+1+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}.$ (IV.77)
By (IV.75), there exists $\varepsilon_{0}>0$ such that for any $\varepsilon\in(0,\varepsilon_{0})$, we have on the one hand that $F_{\varepsilon}\in C^{\infty}(\overline{U},M)$ is well-defined with $F_{\varepsilon}(\overline{U})\subset W^{\prime}$. On the other hand for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, there holds $B_{\varepsilon}(z^{\varepsilon})\subset\phi_{j}(V_{j}\cap\overline{U})$, and there exists $z^{\prime}\in B_{\varepsilon}(z^{\varepsilon})$ satisfying $f\circ\phi_{j}^{-1}$ is differentiable at $z^{\prime}$ with the following inequalities:
@@
    $\big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\leq(2\alpha)^{-1}C_{j}.$ (IV.78)
@@
    $\begin{aligned} &\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z)\big|+\big|(f\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq(2\alpha)^{-1}C_{j}.\end{aligned}$ (IV.79)
@@
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)\big|\\[5.69054pt] &\geq\ \big|d(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|-\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|.\end{aligned}$ (IV.80)